how can I find how many solutions are in an equation like this one?
0=x^2-5x+6

Answer by checkley71(8403)

THERE SHOULD BE 2 SOLUIONS AS THE UNKNOWN IS A FACTOR OF ^2
0=X^2-5X+6
0=(X-6)(X+1)
0=X-6
X=6 SOLUTION
0=X+1
X=-1 SOLUTION